OpenStudy (anonymous):

Please check my answer? Micah paid $12 for 100 photos from the grocery store. Roland paid $18 for 120 photos from the drug store. 1. Write the ratios of the price paid per number of photos purchased for each store. 2. Do the ratios form a proportion? Explain. 1. Write the ratios of the price paid per number of photos purchased for each store. 100/12 = 25/3 120/18 = 20/3 2. Do the ratios form a proportion? Explain. No, the ratios do not form a proportion. To test this, I cross multiplied the two simplified ratios and the two sides were not equal to one another.

OpenStudy (anonymous):

25/3 = 20/3 25*3 = 20*3 75 = 60 This false statement that proves that the ratios are not equal.

OpenStudy (jdoe0001):

sounds good to me

OpenStudy (anonymous):

Thanks!

OpenStudy (jdoe0001):

yw